OpenStudy (anonymous):

HOW DO YOU FIND 20 PERCENT OF A NUMBER

OpenStudy (unklerhaukus):

20% = 20/100 = 20/(5*20) = 1/5 n x 20% = n/5

OpenStudy (anonymous):

.20?

OpenStudy (anonymous):

TIMES . 20?

OpenStudy (unklerhaukus):

that's another way 20% = 20/100 = 0.2 n x 20% = n x 0.2

OpenStudy (unklerhaukus):

(either way will give the same result)